# ImLPR: Image-based LiDAR Place Recognition using Vision Foundation Models

### v1

@@ 4.5 BEV and RIV Representations with Patch-InfoNCE Loss
ImLPR (BEV) 0.945 0.977 0.923 0.968 0.901 0.966 0.869 0.942 0.911 0.965
ImLPR (RIV) 0.979 0.990 0.985 0.993 0.945 0.970 0.844 0.920 0.938 0.968
ImLPR (RIV) 0.990 0.996 0.989 0.995 0.942 0.973 0.888 0.948 0.952 0.978
@@ 6 Limitation
While ImLPR achieves state-of-the-art results on a variety of datasets, its application is limited to homogeneous LiDAR place recognition and is not suited for the emerging area of heterogeneous LiDAR place recognition. The accumulation of a batch of scans, as used in our experiments, partially addresses this issue, but significant differences in field of view of the different LiDAR sensors prevent it from being a complete solution. Furthermore, although we fine-tuned the DINOv2 vision foundation model with MultiConv adapters to better align with the LiDAR domain, addressing heterogeneous LiDAR place recognition or more general problems requires training on larger and more diverse datasets. Despite these constraints, ImLPR’s VFM-based approach advances beyond traditional LPR methods by adapting pre-trained DINOv2 features for LiDAR RIV inputs. This VFM approach enables future development of generalizable place recognition systems through integration with other foundation models, such as those for segmentation and multi-modal processing.
@@ Appendix A Experimental Setup
[…]
@@
Appendix B Data Augmentation in ImLPR
@@ Appendix C Datasets
C.1 HeLiPR Dataset
The HeLiPR dataset comprises six distinct environments—Roundabout01-03, Town01-03, Bridge01-04, DCC04-06, KAIST04-06, and Riverside04-06—captured using four LiDAR sensors: Ouster OS2-128, Velodyne VLP-16C, Livox Avia, and Aeva Aeries II. Each sequence covers approximately 8.5 km, providing sufficient coverage for identifying multiple place recognition pairs. For training, we employ the Ouster OS2-128 sensor across the DCC04-06, KAIST04-06, and Riverside04-06 sequences, yielding a total of 16,435 scans. The test set includes the Roundabout01-03 sequences from both the Ouster OS2-128 and Velodyne VLP-16C sensors, and the Town01-03 sequences from only the Ouster OS2-128 sensor, with a total of 15,375 scans for the Ouster OS2-128 and 7,728 scans for the Velodyne VLP-16C. Sequences are labeled as Sequence-Sensor (e.g., Roundabout-O01 for Ouster, Roundabout-V01 for Velodyne) to distinguish sensor-specific data.
@@ C.3 NCLT Dataset
Appendix D Evaluation on Generalization Capability for Sparse LiDAR Data
Following the intra-session place recognition experiments detailed in the main paper, we further assess the generalization performance of the trained model on the NCLT and HeLiPR-V datasets. Consistent with the primary evaluation, we conduct intra-session place recognition using the HeLiPR-O trained model without additional fine-tuning. Scan accumulation follows the methodology outlined in §C. To ensure a fair comparison, all methods are assessed using the same accumulated scans.
D.1 Evaluation on NCLT Dataset
@@
NCLT (HDL-32E) HeLiPR Roundabout-V (VLP-16C) Total Avg
@@
D.2 Evaluation on HeLiPR-V Dataset
ImLPR surpasses competing methods across nearly all of our evaluation metrics, securing the highest Recall@1 and F1 score for the Roundabout-V01 and Roundabout-V02 sequences, and delivering competitive performance in Roundabout-V03, as detailed in Table. 6. On average, ImLPR achieves a performance improvement more than $10\%$ better than the second-ranked method, BEVPlace++, in both Recall@1 and F1 score. In contrast, other methods show significant performance variability. This is primarily due to the different point cloud distributions which were unobserved during training, which greatly hinder their generalization. Notably, the second-highest F1 scores across the three HeLiPR-V sequences are inconsistent, each attained by a different method, highlighting the difficulty that existing methods have in achieving reliable performance. The Precision-Recall curves presented in Fig. 10 indicate that for Roundabout-V03, ImLPR and MinkLoc3Dv2 yield comparable AUC, followed by CASSPR and BEVPlace++. For Roundabout-V01, however, ImLPR markedly outperforms other methods, with CASSPR, MinkLoc3Dv2, and BEVPlace++ trailing in that order. These results underscore ImLPR’s consistently high performance across the HeLiPR-V dataset.
@@ D.2 Evaluation on HeLiPR-V Dataset
Appendix E Robustness to Yaw Variation
E.1 Place Recognition Performance under Yaw Variation
@@
For a RIV image, yaw variation is equivalent to a horizontal shift. DINOv2, the vision transformer we use, processes patch features via attention mechanisms that capture global inter-patch relationships. Since horizontal shifts reorder patches without modifying their content, the attention mechanism yields nearly translation-equivariant feature representations, preserving the relational structure. Consequently, vision transformers produce similar patch features that are shifted from the original patch features, alongside consistent global tokens for both original and shifted images. Similarly, the MultiConv adapter employs 2D convolutions, which are inherently translation-equivariant, ensuring that a horizontal shift in the input image results in a corresponding shift in patch features while maintaining their values. These refined patch features are aggregated using SALAD’s optimal transport approach, which is invariant to horizontal shifts, as demonstrated in §E.2. Data augmentation strategies, including masking and random transformations, further bolster this stability. However, positional encoding introduces a minor constraint: as yaw rotation shifts the image, identical patches receive different positional encodings, leading to subtle differences in the vision transformer’s output before and after rotation. Despite this, ImLPR achieves more consistent performance than BEVPlace++ under yaw variation, as shown in Table. 7. This highlights the robustness of ImLPR’s model architecture and RIV representation in handling yaw variations effectively.
E.2 Horizontal Shift Invariance in SALAD’s Optimal Transport Aggregation
@@
Appendix F Comparison with 3D Foundation Model

### v2

@@ 4.4 Ablation Studies
ExpB-1 BEV $\mathcal{L}_{\text{TSAP}}$ 0.945 0.977 0.923 0.968 0.901 0.966 0.869 0.942 0.911 0.965
ExpB-2 RIV $\mathcal{L}_{\text{TSAP}}$ 0.979 0.990 0.985 0.993 0.945 0.970 0.844 0.920 0.938 0.968
ExpB-3 RIV $\mathcal{L}_{P},\mathcal{L}_{\text{TSAP}}$ 0.850 0.925 0.833 0.913 0.850 0.943 0.645 0.809 0.795 0.898
ExpB-4 RIV $\mathcal{L}_{P},\mathcal{L}_{\text{TSAP}}$ 0.990 0.996 0.989 0.995 0.942 0.973 0.888 0.948 0.952 0.978
@@ 6 Limitation
While ImLPR achieves SOTA results on a variety of datasets, its application remains limited to homogeneous LPR, where the same type of LiDAR sensor is used for both query and database. It is not yet suited for the emerging task of heterogeneous LPR [34], which requires matching across different LiDAR sensor types with varying specifications and fields of view. Accumulating a batch of scans, as adopted in our experiments, partially mitigates this limitation by improving robustness across sessions, but significant discrepancies in sensor geometry prevent it from serving as a complete solution. Furthermore, although we fine-tuned the DINOv2 vision foundation model with MultiConv adapters to better align with the LiDAR domain, effectively addressing heterogeneous LPR or broader generalization still demands training on larger and more diverse datasets. Despite these constraints, ImLPR’s VFM-based approach extends beyond traditional place recognition methods by adapting pre-trained DINOv2 features to LiDAR RIV inputs. This VFM-based approach enables future development of generalizable place recognition systems through integration with other foundation models, such as those for segmentation [11] and multi-modal processing [13].
ACKNOWLEDGMENT
This project has been partly funded by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00461409), the UKRI project Mobile Robotic Inspector (EP/Z531212/1), and a Royal Society Univ. Research Fellowship (Fallon).
@@ Appendix A Experimental Setup
$\text{F1 score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}},\quad\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}},\quad\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (5)
@@
Appendix B Truncated SmoothAP Loss with Global Image Descriptor
To train ImLPR with the global descriptor, we adopt the Truncated SmoothAP (TSAP) loss [4]. Unlike the triplet and contrastive loss, TSAP optimizes Average Precision (AP) through a continuous approximation, focusing ranking evaluation on top candidates to reduce computation. For a query descriptor $g_{q}$ with positives in $\mathcal{P}^{+}$ and all batch descriptors in $\mathcal{D}$, TSAP is formulated as:
Appendix C Data Augmentation in ImLPR
@@ Appendix D Datasets
D.1 HeLiPR Dataset
The HeLiPR dataset comprises six distinct environments—Roundabout01-03, Town01-03, Bridge01-04, DCC04-06, KAIST04-06, and Riverside04-06—captured using four LiDAR sensors: Ouster OS2-128, Velodyne VLP-16C, Livox Avia, and Aeva Aeries II. Each sequence covers approximately 8.5 km, providing sufficient coverage for identifying multiple place recognition pairs. For training, we employ the Ouster OS2-128 sensor across the DCC04-06, KAIST04-06, and Riverside04-06 sequences, yielding a total of 16,435 scans. The test set includes the Roundabout01-03 sequences from both the Ouster OS2-128 and Velodyne VLP-16C sensors, and the Town01-03 sequences from only the Ouster OS2-128 sensor, with a total of 15,375 scans for the Ouster OS2-128 and 7,728 scans for the Velodyne VLP-16C. Sequences are labeled as Sequence-Sensor (e.g., Roundabout01-O for Ouster, Roundabout01-V for Velodyne) to distinguish sensor-specific data.
@@ D.3 NCLT Dataset
Figure 9: RIV images from HeLiPR (Left) and MulRan (Right) DCC sequence, taken 0.3m apart, vary in appearance: occluded area (sensor location), hidden buildings (range), warped proportions (FoV and sensor location), and shifted colors (intensity distributions).
Table 7: Results in MulRan dataset
MulRan DCC (OS1-64)
Method 01 02 03 Average
R@1 F1 R@1 F1 R@1 F1 AR@1 AF1
LoGG3D-Net (D) 0.117 0.310 0.075 0.144 0.090 0.187 0.094 0.214
CASSPR (D) 0.813 0.943 0.607 0.793 0.629 0.818 0.683 0.851
MinkLoc3dv2 (D) 0.849 0.939 0.651 0.829 0.666 0.841 0.722 0.870
MinkLoc3Dv2 (R) 0.853 0.932 0.632 0.814 0.688 0.832 0.724 0.859
BEVPlace++ (D) 0.811 0.937 0.585 0.766 0.639 0.813 0.678 0.839
BEVPlace++ (R) 0.822 0.922 0.588 0.773 0.655 0.811 0.688 0.835
ImLPR (D) 0.909 0.965 0.861 0.945 0.825 0.918 0.865 0.943
ImLPR (R) 0.918 0.968 0.844 0.940 0.825 0.925 0.862 0.944
Appendix E Evaluation on Generalization Capability
Following the summarized generalization experiments in the main paper, here we further assess the performance of the trained model in detail using three datasets. Consistent with the primary evaluation, we conduct intra-session place recognition using the HeLiPR-O trained model without additional fine-tuning. Scan accumulation follows the methodology outlined in §D. To ensure a fair comparison, all methods are assessed using the same accumulated scans.
E.1 Evaluation on MulRan Dataset
We evaluate the generalization ability of ImLPR on the MulRan DCC01-03 sequences. Although the training and test sets have spatial overlap, they differ substantially due to a four-year temporal gap, differences in LiDAR hardware (OS1-64), and environmental changes such as occlusions. These differences are illustrated in Fig. 9. For evaluation, LiDAR scans are projected into $1022\times 64$ RIV images and resized to $1022\times 126$ to normalize vertical ray differences.
To ensure the robustness of our evaluation protocol and rule out the possibility of performance inflation due to overlap, we additionally replace the original training sequences (DCC04-06) with a geographically seperate sequence, Roundabout04-06, from the HeLiPR with Ouster. We retrain three representative methods, MinkLoc3Dv2 (point-based), BEVPlace++ (BEV-based), and ImLPR (RIV-based), on both training sets: the original (denoted as D) and the revised one (denoted as R).
As shown in Table. 7 and Fig. 10, all methods exhibit only minor differences in performance between the two training sets, validating that the evaluation results are not overly dependent on spatial overlap. Moreover, all models experience performance degradation on the MulRan dataset compared to high-resolution training data, highlighting the impact of domain shift. Among them, ImLPR consistently achieves the highest performance, regardless of training configuration. These results demonstrate that ImLPR generalizes robustly across datasets with different LiDAR sensors and environmental conditions. MinkLoc3Dv2 benefits from sparse convolution that avoids feature generation in occluded regions, ranking second. BEVPlace++ ranks lower, as BEV images tend to retain features even in empty or occluded areas. ImLPR, by contrast, effectively handles low resolution and occlusions through its robust RIV representation and powerful feature extractor. These results demonstrate that ImLPR exhibits strong generalization on the MulRan dataset, even under substantial domain shifts.
E.2 Evaluation on NCLT Dataset
@@
NCLT (HDL-32E) HeLiPR Roundabout-V (VLP-16C) Average
@@
E.3 Evaluation on HeLiPR-V Dataset
ImLPR surpasses competing methods across nearly all of our evaluation metrics, securing the highest Recall@1 and F1 score for the Roundabout01-V and Roundabout02-V sequences, and delivering competitive performance in Roundabout03-V, as detailed in Table. 8. On average, ImLPR achieves a performance improvement more than $10\%$ better than the second-ranked method, BEVPlace++, in both Recall@1 and F1 score. In contrast, other methods show significant performance variability. This is primarily due to the different point cloud distributions which were unobserved during training, which greatly hinder their generalization. Notably, the second-highest F1 scores across the three HeLiPR-V sequences are inconsistent, each attained by a different method, highlighting the difficulty that existing methods have in achieving reliable performance. The Precision-Recall curves presented in Fig. 11 indicate that for Roundabout03-V, ImLPR and MinkLoc3Dv2 yield comparable AUC, followed by CASSPR and BEVPlace++. For Roundabout01-V, however, ImLPR markedly outperforms other methods, with CASSPR, MinkLoc3Dv2, and BEVPlace++ trailing in that order. These results underscore ImLPR’s consistently high performance across the HeLiPR-V dataset.
@@ E.3 Evaluation on HeLiPR-V Dataset
Table 9: Intra-session PR with different threshold in Roundabout01-O
Threshold 1m Threshold 3m Threshold 5m Threshold 7m Threshold 10m
Method R@1 F1 R@1 F1 R@1 F1 R@1 F1 R@1 F1
MinkLoc3Dv2 0.732 0.963 0.933 0.968 0.798 0.888 0.839 0.915 0.933 0.968
BEVPlace++ 0.913 0.959 0.904 0.951 0.909 0.953 0.921 0.959 0.975 0.988
ImLPR 0.965 0.987 0.933 0.970 0.975 0.988 0.960 0.980 0.992 0.996
Appendix F Performance Variations According to Different Thresholds
To rigorously evaluate spatial precision in place recognition, we analyze the impact of varying the distance threshold used to define positive correspondences. While a default threshold of $10\text{\,}\mathrm{m}$, which roughly corresponds to the width of a four-lane road, is a reasonable setting, this choice often results in uniformly high scores across methods. limiting discriminative insight. We therefore systematically tighten the thresholds, and evaluate intra-session performance on Roundabout01-Ouster using three representative methods: MinkLoc3Dv2 (3D point cloud-based), BEVPlace++ (BEV image-based), and ImLPR (RIV image-based).
As presented in Table. 9, ImLPR exhibits consistently superior performance across all threshold levels. Notably, under the most stringent $1\text{\,}\mathrm{m}$ setting, it achieves an R@1 of 0.965 and F1 score of 0.987, substantially outperforming both BEVPlace++ and MinkLoc3Dv2. This indicates that ImLPR is capable of not only recognizing the correct place but also localizing it with high spatial accuracy. BEVPlace++ maintains relatively stable performance across thresholds but shows limited gains under stricter criteria, highlighting its reduced sensitivity to fine-grained spatial alignment. In contrast, MinkLoc3Dv2 exhibits fluctuations due to the sparsity of true positives at lower thresholds, which can degrade metric stability in absence of sufficient positive samples. These results reinforce the robustness of ImLPR under varying spatial tolerances, confirming its capacity for precise place recognition.
Appendix G Robustness to Yaw Variation
G.1 Place Recognition Performance under Yaw Variation
@@
For a RIV image, yaw variation is equivalent to a horizontal shift. DINOv2, the vision transformer we use, processes patch features via attention mechanisms that capture global inter-patch relationships. Since horizontal shifts reorder patches without modifying their content, the attention mechanism yields nearly translation-equivariant feature representations, preserving the relational structure. Consequently, vision transformers produce similar patch features that are shifted from the original patch features, alongside consistent global tokens for both original and shifted images. Similarly, the MultiConv adapter employs 2D convolutions, which are inherently translation-equivariant, ensuring that a horizontal shift in the input image results in a corresponding shift in patch features while maintaining their values. These refined patch features are aggregated using SALAD’s optimal transport approach, which is invariant to horizontal shifts, as demonstrated in §G.2. Data augmentation strategies, including masking and random transformations, further bolster this stability. However, positional encoding introduces a minor constraint: as yaw rotation shifts the image, identical patches receive different positional encodings, leading to subtle differences in the vision transformer’s output before and after rotation. Despite this, ImLPR achieves more consistent performance than BEVPlace++ under yaw variation, as shown in Table. 10. This highlights the robustness of ImLPR’s model architecture and RIV representation in handling yaw variations effectively.
G.2 Horizontal Shift Invariance in SALAD’s Optimal Transport Aggregation
@@
Appendix H Ablation Studies
Table 11: Ablation study for adapter and trained block
Method Adapter
Trained
Block
Roundabout-O Town-O DCC Roundabout-V Average
AR@1 AF1 AR@1 AF1 AR@1 AF1 AR@1 AF1 AR@1 AF1
ExpC-1 0 0.321 0.530 0.379 0.553 0.376 0.555 0.249 0.464 0.331 0.526
ExpC-2 2 0.850 0.925 0.833 0.913 0.850 0.943 0.645 0.809 0.795 0.898
ExpC-3 0 0.972 0.987 0.965 0.983 0.928 0.965 0.799 0.906 0.916 0.960
ExpC-4 2 0.990 0.996 0.989 0.995 0.942 0.973 0.888 0.948 0.952 0.978
ExpC-5 6 0.990 0.995 0.990 0.996 0.941 0.973 0.868 0.940 0.947 0.976
ExpC-6 10 0.986 0.994 0.988 0.995 0.935 0.972 0.870 0.936 0.945 0.974
H.1 MultiConv Adapter and the Number of Trained Block in DINOv2
VFMs, such as DINOv2, leverage extensive pre-training on large-scale datasets to deliver robust feature representations. To preserve these learned representations during adaptation for LPR, it is essential to avoid catastrophic forgetting while effectively bridging the domain gap between natural vision images and RIV images. Fine-tuning multiple transformer blocks risks overwriting the model’s general knowledge, whereas minimal fine-tuning with strategic adaptations can maintain performance. Following the main evaluation, this section examines how the inclusion of the MultiConv adapter and varying the number of trained transformer blocks influence the performance of ImLPR.
The results, presented in Table. 11, highlight the critical role of the MultiConv adapter. Without the adapter and without fine-tuning (ExpC-1), the model fails to perform effective place recognition due to the significant domain gap between natural images and RIV images. Fine-tuning without the adapter (ExpC-2) also results in reduced performance, as the limited number of trainable parameters hinders the model’s ability to extract robust features from RIV images. In contrast, incorporating the MultiConv adapter while keeping most transformer blocks frozen (ExpC-3 and ExpC-4) significantly improves performance. This demonstrates the adapter’s ability to efficiently address domain shifts, leveraging DINOv2’s pre-trained knowledge for enhanced place recognition.
Table 12: Ablation study for yaw change and mask types
Method
Yaw
Change
Line & Square
Mask
Cylindrical
Mask
Roundabout-O Town-O DCC Roundabout-V Average
AR@1 AF1 AR@1 AF1 AR@1 AF1 AR@1 AF1 AR@1 AF1
ExpD-1 0.937 0.972 0.943 0.973 0.905 0.966 0.781 0.893 0.892 0.951
ExpD-2 0.980 0.990 0.958 0.980 0.896 0.983 0.821 0.927 0.914 0.970
ExpD-3 0.984 0.992 0.975 0.989 0.894 0.985 0.837 0.928 0.923 0.974
ExpD-4 0.990 0.996 0.989 0.995 0.942 0.973 0.888 0.948 0.952 0.978
H.2 Effect of Data Augmentation
We evaluate the impact of data augmentation on ImLPR’s inter-session place recognition performance, with results presented in Table. 12. Without any augmentations (ExpD-1), the model achieves baseline performance but struggles with orientation variability and projection artifacts in RIV images, limiting its robustness. Applying random yaw rotation (ExpD-2) significantly improves performance by introducing column shifts that enhance the model’s resilience to orientation changes in LiDAR scans. Further improvement is observed with the addition of line and square masking (ExpD-3). Both masks simulate sparse regions by introducing empty pixels in RIV images, mimicking the effect of dropout and training the model to perform place recognition with partially empty patch, thus boosting robustness to incomplete inputs.
The best performance is achieved when combining all augmentations—random yaw rotation, line, square, and cylindrical masking (ExpD-4). The cylindrical mask also enhances the model’s ability to handle the extreme occlusions by training it to rely on partial RIV images. This configuration delivers superior results across all datasets, with notable gains on DCC and Roundabout-V, which include occlusions in their scans. These results underscore the vital role of data augmentation in enhancing ImLPR’s robustness, with the combined yaw rotation and masking strategies effectively addressing orientation variability, projection artifacts, and scene sparsity in LPR.
H.3 Dimension of DINOv2
The impact of varying the DINOv2 backbone’s feature dimension on ImLPR’s inter-session place recognition performance and computational efficiency is assessed in Table. 13. All experiments here use a single NVIDIA GeForce RTX 3090 GPU to measure computation time. This ablation study evaluates three DINOv2 configurations—ViT-S/14, ViT-B/14, and ViT-L/14—with feature dimensions of 384, 768, and 1024, respectively, analyzing their place recognition efficacy and computational cost. The final two transformer blocks are fine-tuned for all models to ensure consistent adaptation, and computational efficiency is compared against other SOTA methods.
H.3.1 Performance Analysis
The place recognition performance of ImLPR across different feature dimensions is detailed in Table. 13, which presents results for multiple datasets. All three configurations—ViT-S/14 (ExpE-1), ViT-B/14 (ExpE-2), and ViT-L/14 (ExpE-3)—exhibit similar performance, achieving robust place recognition outcomes. This similarity indicates that higher feature dimensions do not always yield proportional improvements in accuracy. Notably, the smallest model, ViT-S/14, demonstrates exceptional robustness, particularly on Roundabout-O, Town-O, and Roundabout-V, where it delivers highly competitive results. These findings indicate that larger models may not always outperform smaller ones, as their increased complexity can hinder effective training on diverse RIV images and lead to overfitting on the training data.
Table 13: Ablation study for feature dimension and computation time
Method
Feature
Dim.
Runtime
Roundabout-O Town-O DCC Roundabout-V Average
AR@1 AF1 AR@1 AF1 AR@1 AF1 AR@1 AF1 AR@1 AF1
ExpE-1 384 18.1 0.990 0.996 0.989 0.995 0.942 0.973 0.888 0.948 0.952 0.978
ExpE-2 768 23.3 0.989 0.995 0.988 0.994 0.943 0.973 0.861 0.942 0.945 0.976
ExpE-3 1024 58.0 0.989 0.995 0.986 0.994 0.957 0.979 0.873 0.944 0.951 0.978
H.3.2 Computational Cost
The computational efficiency is critical for real-world place recognition applications. As reported in Table. 13, the time required for descriptor extraction scales with feature dimension: ViT-S/14 (ExpE-1) is the most efficient, followed by ViT-B/14 (ExpE-2), with ViT-L/14 (ExpE-3) incurring the highest computational cost. All configurations achieve extraction times below 100ms, demonstrating their suitability for real-time performance in LPR. As depicted in Fig. 13, we further compare ImLPR against baseline methods using Average Recall@1 on Roundabout-O and Town-O. BEVPlace++ requires 27.5ms for descriptor extraction, despite a smaller feature dimension of 128, surpassing the runtime of ViT-S/14 and ViT-B/14, due to its reliance on multiple ResNet instances for yaw invariance. MinkLoc3Dv2, utilizing 3D sparse convolution, achieves the fastest extraction time but yields the lowest Average Recall@1. In contrast, CASSPR’s attention-based networks and per-point feature extraction lead to the longest computation time. ImLPR’s Vision Transformer-based backbone inherently supports robustness to yaw variation and achieve a low latency of 18.1ms. Furthermore, as shown in Fig. 13, ImLPR achieves the highest Average Recall@1, demonstrating its suitability for real-time place recognition applications requiring minimal latency.
Table 14: Average Results of Intra-session PR with Visual Place Recognition models
DINOv2 Runtime Roundabout-O Town-O Average
Method Backbone (ms) AR@1 AF1 AR@1 AF1 AR@1 AF1
SALAD ViT-B/14 21.3 0.841 0.919 0.771 0.874 0.806 0.896
BoQ ViT-B/14 22.7 0.934 0.967 0.896 0.947 0.915 0.957
SelaVPR (global) ViT-L/14 79.5 0.955 0.978 0.941 0.970 0.948 0.974
ImLPR ViT-S/14 18.1 0.986 0.994 0.966 0.984 0.976 0.989
Appendix I Comparison with Visual Place Recognition Models
ImLPR employs a VFM and adopts image-based aggregation strategies, SALAD, similar to those used in VPR. To assess whether RIV images—composed of reflectivity, range, and normal ratio channels—can be effectively processed by conventional VPR models, we compare ImLPR with recent VPR methods, including SALAD [8], BoQ [36], and SelaVPR [37]. To ensure fair comparison, all models were trained using the same RIV inputs and identical data augmentation strategies. We also used the same loss function (TSAP loss) across all methods, and all feature encoders and training hyperparameters were kept at their default settings.
Table. 14 reports average intra-session place recognition performance on Roundabout-O and Town-O. While existing VPR methods achieve high accuracy in traditional visual tasks, they show limited performance when directly applied to RIV images. SALAD and BoQ, which use ViT-B/14 backbones trained with the TSAP loss, underperform compared to ImLPR across all metrics, indicating that naive application of VPR pipelines to LiDAR-based images does not adequately handle the sensor domain gap. SelaVPR, which integrates adapters and uses a larger ViT-L/14 backbone, performs better than SALAD and BoQ but still falls short of ImLPR.
Despite using a smaller backbone (ViT-S/14), ImLPR outperforms all baselines by a notable margin and operates with the lowest runtime (18.1 ms). This highlights both the computational efficiency and accuracy of the proposed method. The performance gap emphasizes the importance of explicitly addressing the domain shift between camera and LiDAR data. In particular, the use of adapters allows for better alignment with LiDAR-specific structures, as evidenced in both SelaVPR and ImLPR, while our Patch-InfoNCE loss further enhances RIV feature learning by leveraging range-aware supervision, which is unavailable in image-based models. These results demonstrate that high performance on LiDAR images cannot be achieved by merely applying VPR models composed of DINOv2 and an aggregation network to RIV inputs. Instead, effective place recognition requires proper adaptation beyond simple fine-tuning, as well as LiDAR-aware feature learning.
Appendix J Comparison with 3D Foundation Model